hello,
Please help me solve this problem. I do not know to approach it.
Solve with quadratic formula
3/4(x^2)-2x+1/2=0

Answer by TimothyLamb(4379)

(3/4)x^2 - 2x + (1/2) = 0
3/4 = 0.75
1/2 = 0.5
0.75x^2 - 2x + 0.5 = 0
---
the above quadratic equation is in standard form, with a=0.75, b=-2, and c=0.5
---
to solve the quadratic equation, by using the quadratic formula, plug this:
0.75 -2 0.5
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the two roots (i.e. the two solutions), of the quadratic are:
x = 2.38742589
x = 0.27924078
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php